OpenStudy (anonymous):

Use your knowledge of the derivative to compute the value of the limit given below: \lim_{h \rightarrow 0} \frac{\sqrt[1]{(x + h)^{3}} - \sqrt[1]{x^{3}}}{h} . The derivative is \frac{dy}{dx}= .

OpenStudy (anonymous):

\[\lim_{h \rightarrow 0} \frac{\sqrt[1]{(x + h)^{3}} - \sqrt[1]{x^{3}}}{h}\]

OpenStudy (anonymous):

1st root?

OpenStudy (dumbcow):

square root always has index of 2...index of 1 means radical not needed i assumed it was a typo

OpenStudy (anonymous):

thx dc.

OpenStudy (dumbcow):

find derivative of x^(3/2)

OpenStudy (anonymous):

this is the limit defition for the derivivative of x^(3/2)

OpenStudy (anonymous):

none of these work :(

OpenStudy (dumbcow):

@rukh is the question written with a 1 as index of the radical?